OpenStudy (anonymous):

What is the equation of the line passing through (–6, 1) and perpendicular to the line y = –3x + 1 in slope-intercept form?

OpenStudy (anonymous):

i need big help on this ASAP please

OpenStudy (ajprincess):

Slope of the given line*Slope of the required line=-1 -3*m=-1 m=1/3  Can u find nw?